OpenStudy (anonymous):

Find two consecutive integers whose sum is 113. Which of the following equations could be used to solve the problem? 2x= 113 2x+ 1 = 113 2x+ 2 = 113 x2+ 1= 113

OpenStudy (anonymous):

If x is a number. x+1 is the next number. If you add them to get 113, then you have equation: (x)+(x+1)=113 or: x+x+1=113 OR 2x+1=113

OpenStudy (anonymous):

2x + 1 = 113

OpenStudy (anonymous):

2 numbers are 56 57